ADNAN MUHAMMED

Grade 12,

Simplify and solve the following linear equations. 3(t – 3) = 5(2t + 1)

Simplify and solve the following linear equations. 
3(t – 3) = 5(2t + 1) 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

3(t–3) = 5(2t + 1)
⇒ 3t- 9 = 10t + 5
⇒ 3t- 10t = 5 + 9
⇒-7t = 14
⇒t = 14/-7
⇒t = -2

Thanks